Suppose you’re evaluating three alternative MMMF investments. The first fund buys a diversified portfolio of municipal securities from across the country and yields 2.9 percent. The second fund buys only taxable, short-term commercial paper and yields 5.4 percent. The third fund specializes in the municipal debt from the state of New Jersey and yields 3.4 percent. Your federal income tax rate is 35 percent and you are a resident of Texas, which has no state income tax.

a. Calculate the after-tax yield for each of the alternatives. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

After tax yield

Municipal fund ?   %

taxable fund ?   %

New Jersey municipal fund ?   %
Answer #1

After-tax Yield = Yield * [1 - (Federal tax rate + State Tax rate)]

Municipal Fund = 2.9% * [1 - (0.0 + 0.0)] = 2.90%

Taxable Fund = 5.4% * [1 - (0.35 + 0.0)] = 3.51%

New Jersey Municipal Fund = 3.4% * [1 - (0.0 + 0.0)] = 3.40%